Solve each system of equations by graphing:
(y = x +3
(y = - 2x - 3

Answer by mananth(16949)

(y = x +3
(y = - 2x - 3

Find the x & y intercept points by plugging x=0,y=0 in both equations.
equation I y = x +3
when x= 0 y= 3 ,( 0 , 3 )
when y = 0 x= 3 ,( -3 , 0 )
Equation 2 y = - 2x - 3

when x= 0 y= -3 ( 0 , -3 )
when y = 0 x= -1.5 ( -1.5 , 0 )



You will observe that the point of intersection is -2 , 1
The solution of the system is ( -2 , 1 )